You are analyzing the value of a potential unicorn your Investment Banker is trying to sell you on. He says that its worth today is $1 billion. While you are comfortable with the relative accuracy of the near term cash flows, you are interested in figuring out how much he is expecting the unicorn to be worth at the end of year 4. Assume a 15% discount rate. You expect its cash flows over the next 4 years to be as shown below.

Year: Cash Flow

0. -20M

1. -10M

2 0

3 12M

4 40M

What is this unicorn’s future value at the end of year 4? * A). $ 2,064,672 B). $997,064,672 C). $1,475,395,126 D). $1,695,553,239 E). $1,756,892,730
Answer #1

Let the future value at the end of year 4 be V

The present value of the Unicorn ($ 1 billion) is equal to the sum of present value of all cashflows till year 4 plus the present value of Unicorn at the end of year 4. Putting all the figures in millions of dollars and since 1 billion = 1000 million

1000 = -20 - 10/1.15 + 0/1.15^2 + 12/1.15^3+40/1.15^4 + V/1.15^4

=> 1000 = 2.064672+V/1.15^4

'=> V = 997.9353 * 1.15^4

=1745.395125 million

= $1,745,395,125

( Please check the options again)